# TacDiffusion: Force-domain Diffusion Policy for Precise Tactile Manipulation

### v1

@@ I Introduction
Throughout the history of robotics research, the importance of tactile feedback and force control for high-precision assembly has been consistently recognized [7, 8, 9, 10, 11, 12, 5, 13]. However, several challenges persist in precise force control, including the difficulty of accessing to appropriate robot hardware and expensive force sensors, the complexity of ensuring stability and safety while regulating force, the sensitivity of force control to environmental changes, the difficulty of estimating environment constraints and contact dynamics in dynamic settings, and the challenge of collecting high-quality tactile data for learning force control. Due to these barriers, the use of simpler motion-domain action spaces, with impedance control as an indirect force control method is often favored by the robot learning community. Nevertheless, the increasing diversity of contact-rich manipulation tasks highlights the equal importance of simultaneously regulating motion, compliance, and force, so that agents can autonomously perform a wide range of task stably and robustly, without the need for explicit controller switching [14]. Despite the recent successes of implementing transformer [15, 16, 17, 18, 19, 20] and/or diffusion-based [21, 22, 23, 24, 25, 26, 27, 28, 29, 30] policies for robot manipulation that exhibit excellent generalization capability, it remains unexplored how to integrate force control with these models for high-precision tactile manipulation, so that the benefits of these generative models for multi-modal modelling and prediction can be fully exploited.
@@ II-B Enhancing Transferability in Robotic Assembly
In the last decade, there is extensive literature on generating robotic assembly policies with broad generalization. Deep Reinforcement Learning-based methods, for instance, typically achieve the generalization ability through training with multiple objects [32, 33]. Another noteworthy case is meta-learning, which trains a pre-trained model using online or offline data from a diverse and comprehensive set of tasks, enabling domain adaptation ability through fine-tuning [34, 35]. Furthermore, sim-to-real based approaches have gained attention for their cost-effective data collection in the simulation environments, and zero-shot sim-to-real transfer for perception-initialized assembly has been only recently demonstrated [36, 37]. Besides, to tackle precise manipulation, RVT-2 [20] trained a transformer-based multi-task policy. Despite improving performance on multi-task learning benchmark, its success rate on high-precision (millimeter level) insertion tasks, roughly 50%, is far from being satisfactory to deploy to real assembly production. Aside from these approaches, evolutionary algorithms with parameterized robot skills have shown transferability across tasks via fine-tuning [31]. However, achieving zero-shot transfer on high-precision tasks with a satisfactory success rate in the real world remains an open challenge
@@ III-A Framework Overview
During the assembly process, the interaction between the robot and the environment is captured as observation $\bm{o}$, which includes the external wrench, internal wrench, and end-effector’s speed. The diffusion model then predicts the force-domain actions ($\bm{a}:=\bm{F}_{df}$) based on both the current observation $\bm{o}_{curr}$ and the previous observation $\bm{o}_{prev}$. Due to the restrictions of computational resources, the diffusion model’s inference frequency typically ranges from 50 Hz to 500 Hz (Table I), which is misaligned with the robot’s 1000 Hz real-time control loop. To mitigate this, we designed a dynamic system-based filter to interpolate the diffusion model’s predictions $\bm{F}_{df}$. The filtered action is then transmitted to the impedance controller with feed-forward force. Based on the desired goal $\bm{x}_{d}$ (insertion hole’s pose) and the force command, it regulates the robot’s motion and force behavior simultaneously.
@@ III-B1 Diffusion Process
Furthermore, the noise $\bm{\epsilon}_{\tau}$ also plays a crucial role in the subsequent denoising process. To account for this, we construct the noise estimator $\hat{\bm{\epsilon}}(\cdot)$ using a residual neural network [42], as illustrated in Fig. 1, and train it by minimizing the following loss function:
@@ V Conclusion
In this work, we present a novel framework leveraging diffusion models to generate 6D wrench for tactile manipulation in high-precision robotic assembly tasks. Our approach, being the first force-domain diffusion policy, demonstrated excellent improved zero-shot transferability compared to prior work, by achieving an overall 95.7% success rate in zero-shot transfer in experimental evaluations. Additionally, we investigate the trade-off between accuracy and inference speed and provide a practical guideline for optimal model selection. Further, we address the frequency misalignment between the diffusion policy and the real-time control loop with a dynamic system-based filter, significantly improving the task success rate by 9.15%. Extensive experimental studies in our work underscores the effectiveness of our framework in real-world settings, showcasing a promising approach tackling high-precision tactile manipulation by learning diffusion-based transferable skills from expert policies containing primitive-switching logic. In future work, we will focus on extending the framework’s applicability to a broader range of high-precision assembly tasks and integrating additional sensing modalities to enhance system adaptability and robustness in real-time environments.

### v2

@@ I Introduction
Throughout the history of robotics research, the importance of tactile feedback and force control for high-precision assembly has been consistently recognized [7, 8, 9, 10, 11, 12, 5, 13]. However, several challenges persist in precise force control, including the difficulty of accessing to appropriate robot hardware and expensive force sensors, the complexity of ensuring stability and safety while regulating force, the sensitivity of force control to environmental changes, the difficulty of estimating environment constraints and contact dynamics in dynamic settings, and the challenge of collecting high-quality tactile data for learning force control. Due to these barriers, the use of simpler motion-domain action spaces, with impedance control as an indirect force control method is often favored by the robot learning community. Nevertheless, the increasing diversity of contact-rich manipulation tasks highlights the equal importance of simultaneously regulating motion, compliance, and force, so that agents can autonomously perform a wide range of task stably and robustly, without the need for explicit controller switching [14]. Despite the recent successes of implementing transformer [15, 16, 17, 18, 19, 20] and/or diffusion-based [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] policies for robot manipulation that exhibit excellent generalization capability, it remains unexplored how to integrate force control with these models for high-precision tactile manipulation, so that the benefits of these generative models for multi-modal modelling and prediction can be fully exploited.
@@ II-B Enhancing Transferability in Robotic Assembly
In the last decade, there is extensive literature on generating robotic assembly policies with broad generalization. Deep Reinforcement Learning-based methods, for instance, typically achieve the generalization ability through training with multiple objects [33, 34]. Another noteworthy case is meta-learning, which trains a pre-trained model using online or offline data from a diverse and comprehensive set of tasks, enabling domain adaptation ability through fine-tuning [35, 36, 37]. Furthermore, sim-to-real based approaches have gained attention for their cost-effective data collection in the simulation environments, and zero-shot sim-to-real transfer for perception-initialized assembly has been only recently demonstrated [38, 39]. Besides, to tackle precise manipulation, RVT-2 [20] trained a transformer-based multi-task policy. Despite improving performance on multi-task learning benchmark, its success rate on high-precision (millimeter level) insertion tasks, roughly 50%, is far from being satisfactory to deploy to real assembly production. Aside from these approaches, evolutionary algorithms with parameterized robot skills have shown transferability across tasks via fine-tuning [32]. However, achieving zero-shot transfer on high-precision tasks with a satisfactory success rate in the real world remains an open challenge
@@ III-A Framework Overview
During the assembly process, the interaction between the robot and the environment is captured as observation $\bm{o}$, which includes the external wrench, internal wrench, and end-effector’s speed. The diffusion model then predicts the force-domain actions ($\bm{a}:=\bm{F}_{df}$) based on both the current observation $\bm{o}_{curr}$ and the previous observation $\bm{o}_{prev}$. Due to the restrictions of computational resources, the diffusion model’s inference frequency typically ranges from 50 Hz to 500 Hz (Table I), which is misaligned with the robot’s 1000 Hz real-time control loop. To mitigate this, we design a dynamic system-based filter to interpolate the diffusion model’s predictions $\bm{F}_{df}$. The filtered action is then transmitted to the impedance controller with feed-forward force. Based on the desired goal $\bm{x}_{d}$ (insertion hole’s pose) and the force command, it regulates the robot’s motion and force behavior simultaneously.
@@ III-B1 Diffusion Process
Furthermore, the noise $\bm{\epsilon}_{\tau}$ also plays a crucial role in the subsequent denoising process. To account for this, we construct the noise estimator $\hat{\bm{\epsilon}}(\cdot)$ using a residual neural network, as illustrated in Fig. 1, and train it by minimizing the following loss function:
@@ V Conclusion
In this work, we present a novel framework leveraging diffusion models to generate 6D wrench for tactile manipulation in high-precision robotic assembly tasks. Our approach, being the first force-domain diffusion policy, demonstrated excellent improved zero-shot transferability compared to prior work, by achieving an overall 95.7% success rate in zero-shot transfer in experimental evaluations. Additionally, we investigate the trade-off between accuracy and inference speed and provide a practical guideline for optimal model selection. Further, we address the frequency misalignment between the diffusion policy and the real-time control loop with a dynamic system-based filter, significantly improving the task success rate by 9.15%. Extensive experimental studies in our work underscore the effectiveness of our framework in real-world settings, showcasing a promising approach tackling high-precision tactile manipulation by learning diffusion-based transferable skills from expert policies containing primitive-switching logic. In future work, we will focus on extending the framework’s applicability to a broader range of high-precision assembly tasks and integrating additional sensing modalities to enhance system adaptability and robustness in real-time environments.
ACKNOWLEDGMENT
The authors acknowledge the financial support by the Bavarian State Ministry for Economic Affairs, Regional Development and Energy (StMWi) for the Lighthouse Initiative KI.FABRIK (Phase 1: Infrastructure as well as the research and development program under, grant no. DIK0249), the European Union’s Horizon 2020 research and innovation programme as part of the project euROBIN under grant no. 10107059, the Federal Ministry of Education and Research of Germany (BMBF) in the programme of “Souverän. Digital. Vernetzt.” Joint project 6G-life, project identification number 16KISK002, and by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden.